OpenStudy (anonymous):

f(x)=x^2-1; g(t)=sqrt t+1. compute a. f(g(x))

OpenStudy (anonymous):

Let\[f(x)=x^2-1\]and\[g(x)=\sqrt{x+1}.\]Then\[f(g(x))=(\sqrt{x+1})^2-1=x.\]